OpenStudy (anonymous):

Evaluate. Assume u>0 when ln u appears.

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 8 }{ (y-9)^3 }dy\]

OpenStudy (anonymous):

Woops, just realized I was being retarded about this earlier. 8∫(y-9)^-3dy = (-8/2)(y-9)^-2 + c I don't think there's actually any need for lns.